# Task-Agnostic Machine Learning-Assisted Inference

### v1

@@ 3.2 Theoretical guarantees
In this section, we examine the theoretical properties of PSPS. In what follows, $\stackrel{{\scriptstyle P}}{{\rightarrow}}$ denotes convergence in probability and $\stackrel{{\scriptstyle D}}{{\rightarrow}}$ denotes convergence in distribution.
@@
$n^{1/2}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}-{\bm{\theta}}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left(0,{\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})\right),$ (17)
where ${\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})={\bf V}(\widehat{\bm{\theta}}_{\cal L})-{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$.
Assume the $k$-th column of ${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ is not a zero vector and at least one of ${\bf V}(\widehat{\bm{\eta}}_{\cal L})$ and ${\bf V}(\widehat{\bm{\eta}}_{\cal U})$ are positive definite, then ${\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}\leq{\bf V}(\widehat{\bm{\theta}}_{\cal L})_{kk}$.
This result indicates a greater reduction in variance for the ML-assisted estimator is associated with larger values of ${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ and smaller values of ${\bf V}(\widehat{\bm{\eta}}_{\cal L})$, ${\bf V}(\widehat{\bm{\eta}}_{\cal U})$, and $\rho$. The variance reduction term $[{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})]_{kk}$ can also serve as a metric for selecting the optimal ML model in ML-assisted inference.
Our next result shows that three existing methods, i.e., PPI, PPI++, and POP-Inf, are asymptotically equivalent to PSPS with different weighting vector. A broader class for consistent estimator of $\bm{\theta}^{*}$ is $\widehat{{\bm{\theta}}}(\bm{\omega})=\bm{\omega}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\bm{\omega}^{\rm T}\widehat{{\bm{\theta}}}_{{\cal L}}-\widehat{\bm{\eta}}_{{\cal L}})$, where $\bm{\omega}$ is a $K\times K$ matrix. The consistency of $\widehat{{\bm{\theta}}}(\bm{\omega})$ for $\bm{\theta}^{*}$ only requires $\bm{\omega}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}\mathbf{0}$. Since $(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}\mathbf{0}$, assigning arbitrary fixed weights for will satisfy the condition. However, the choice of weights influences the efficiency of the estimator as illustrated in proposition 2 later.
@@
Assuming equation (16) and regularity condition for the asymptotically normality of current ML-assisted estimator holds. For M-estimation problem, we have
$\widehat{{\bm{\theta}}}\left({\rm diag}(\bm{\omega}_{\textnormal{ele}}){\bf C}\right)-\widehat{{\bm{\theta}}}_{\textnormal{{POP-Inf}}}\stackrel{{\scriptstyle D}}{{\rightarrow}}{\bf 0},\quad\widehat{{\bm{\theta}}}\left({\rm diag}(\bm{\omega}_{\textnormal{tr}}){\bf C}\right)-\widehat{{\bm{\theta}}}_{\textnormal{{PPI++}}}\stackrel{{\scriptstyle D}}{{\rightarrow}}{\bf 0},\quad\widehat{{\bm{\theta}}}\left({\rm diag}({\bf 1}){\bf C}\right)-\widehat{{\bm{\theta}}}_{\textnormal{{PPI}}}\stackrel{{\scriptstyle D}}{{\rightarrow}}{\bf 0}.$
@@
This demonstrates that for M-estimation problems, our method is asymptotically equivalent to POP-Inf, PPI++, and PPI each with the respective weights ${\rm diag}(\bm{\omega}_{\textnormal{ele}}){\bf C}$, ${\rm diag}(\bm{\omega}_{\textnormal{tr}}){\bf C}$, and ${\rm diag}({\bf 1}){\bf C}$. Therefore, PSPS can be viewed as a generalization of these existing methods.
@@
Together with proposition 1, our results demonstrate that our protocol provides a more efficient estimator compared to existing methods for the M-estimation problem. Furthermore, the applicability of our protocol is not limited to M-estimation and only requires summary statistics as input. It also indicates that in a distributed inference setting (Jordan et al.,, 2018) where individual-level data are not available, PSPS proves to be the optimal approach for combining shared summary statistics.
@@ 3.3 Extensions
We also provide several extensions to ensure the broad applicability of our proposed method.
@@ 3.3.1 Labeled data and unlabeled data are not independent
Here, we relax the assumption that the labeled data and unlabeled data are independent. When they are not independent, this can leads to the non-zero covariance between the $\widehat{\bm{\eta}}_{\cal L}$ and $\widehat{\bm{\eta}}_{\cal U}$. Consider a more broader class of summary statistics asymptotically satisfying
$n^{1/2}\left(\begin{array}[]{c}\widehat{{\bm{\theta}}}_{\cal L}-{\bm{\theta}}\\ \widehat{\bm{\eta}}_{\cal L}-\bm{\eta}\\ \widehat{\bm{\eta}}_{\cal U}-\bm{\eta}\\ \end{array}\right)\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left\{\left(\begin{array}[]{c}{\bf 0}_{K}\\ {\bf 0}_{K}\\ {\bf 0}_{K}\end{array}\right),\left(\left(\begin{array}[]{ccc}{\bf V}(\widehat{\bm{\theta}}_{\cal L})&{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})\\ {\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L})&\sqrt{\rho}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})\\ {\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})&\sqrt{\rho}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})&\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U})\end{array}\right)\right)\right\}$
We can similarly employ the one-step debiasing
$\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}^{\texttt{ no-indep}}=\bm{\omega}_{0}^{\rm T}\widehat{\bm{\eta}}_{\cal U}-\bm{\omega}_{0}(\widehat{\bm{\eta}}_{\cal L}-\widehat{{\bm{\theta}}})$ (18)
where $\bm{\omega}_{0}=({\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})-{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{\rm T}({\bf V}({\widehat{\bm{\eta}}_{\cal L}})+{\bf V}({\widehat{\bm{\eta}}_{\cal U}})-2{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{-1}$ and $\operatorname{Var}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}^{\texttt{ no-indep}})=\operatorname{Var}(\widehat{\bm{\theta}}_{\cal L})-({\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})-{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{\rm T}({\bf V}({\widehat{\bm{\eta}}_{\cal L}})+{\bf V}({\widehat{\bm{\eta}}_{\cal U}})-2{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{-1}({\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})-{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))$. The theoretical guarantees of the proposed estimator can be similarly derived by theorem 1.
@@ 3.3.2 Sensitivity analysis for distributional shift between labeled and unlabeled data
The other assumption of our approach is that the labeled and unlabeled data are identically distributed so that we can ensure $\widehat{\bm{\eta}}_{\cal L}-\widehat{\bm{\eta}}_{\cal U}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bf 0}$ and validity of PSPS results. To address the potential violation of this assumption, we introduce a sensitivity analysis with hypothesis testing for the null $H_{0}:\bm{\eta}_{{\cal L},k}=\bm{\eta}_{{\cal U},k}$ with test statistics
$\frac{\widehat{\bm{\eta}}_{{\cal L},k}-\widehat{\bm{\eta}}_{{\cal U},k}}{\sqrt{\operatorname{Var}(\widehat{\bm{\eta}}_{{\cal L},k})+\operatorname{Var}(\widehat{\bm{\eta}}_{{\cal U},k})}}\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}(0,1)$ (19)
to assess if $\bm{\eta}_{{\cal L},k}$ and $\bm{\eta}_{{\cal U},k}$ are significantly different. Here, the subscript $k$ indicates the $k$-th coordinate. We recommend using p-value $<0.1$ as evidence for heterogeneity and caution the interpretation for results of ML-assisted inference.
4 ML-assisted False Discovery Rate Control
Controlling FDR is crucial for statistical application in various scientific disciplines (Benjamini and Hochberg,, 1995; Benjamini and Yekutieli,, 2001; Barber and Candès,, 2015). In this section, we show that our protocol can be utilized for ML-assisted FDR control, achieving greater statistical power compared to classical methods that solely rely on labeled data. We consider the linear regression model with
$\bm{Y}=\sum_{k=1}^{K}\bm{X}_{k}\beta_{k}+\bm{\epsilon},\bm{\epsilon}\sim\mathcal{N}\left(\mathbf{0},\sigma^{2}\bm{I}_{n}\right)$ (20)
Here, $\bm{Y}$ is a $n$-dimensional vector, $\bm{X}_{k}$ is also a $n$-dimensional vector representing the $k$-th feature. A central question here is to select the variables associated with the outcome $\bm{Y}$ after accounting for all other variables. Formally, denote $\widehat{S}\subset\{1,2,\ldots,K\}$ as the set of selected variables. The FDR is defined as
$\text{FDR}=\mathbb{E}[\frac{\#\left\{k:\beta_{k}=0,k\in\widehat{S}\right\}}{\#\{k:k\in\widehat{S}\}\vee 1}].$ (21)
We consider two popular FDR control methods: Benjamini-Hochberg (BH) procedure (Benjamini and Hochberg,, 1995) and knockoff (Barber and Candès,, 2015). BH procedure first computes p-values $p_{k}$ for each feature ${\bf X}_{k}$ using multiple linear regression and then applies a threshold to these p-values. Features with p-values smaller than the threshold are considered as discoveries. The knockoff method does not rely on p-values. It includes three key components: a ranking algorithm that assigns an importance metric to each variable, an augmented design that creates fake variables, and a symmetric statistic that utilizes the importance metric for each original variable to control the final FDR (Ke et al.,, 2024).

### v3

@@ 3.2 Theoretical guarantees
In this section, we examine the theoretical properties of PSPS. In what follows, $\stackrel{{\scriptstyle P}}{{\rightarrow}}$ denotes convergence in probability and $\stackrel{{\scriptstyle D}}{{\rightarrow}}$ denotes convergence in distribution. All proofs are deferred to the appendix A.
@@
$n^{1/2}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}-{\bm{\theta}}^{*})\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left({\bf 0},{\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})\right),$
where ${\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})={\bf V}(\widehat{\bm{\theta}}_{\cal L})-{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$. Assume the $k$-th column of ${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ is not a zero vector and at least one of ${\bf V}(\widehat{\bm{\eta}}_{\cal L})$ and ${\bf V}(\widehat{\bm{\eta}}_{\cal U})$ are positive definite, then ${\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})_{kk}\leq{\bf V}(\widehat{\bm{\theta}}_{\cal L})_{kk}$. With $\widehat{\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}),\lim_{n}\mathbb{P}(\theta_{k}^{{}^{*}}\in\mathcal{C}_{\alpha,k}^{\textnormal{{PSPS}}})=1-\alpha$.
$\widehat{\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})$ can be obtained by applying the algebraic form of ${\bf V}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}})$ using the bootstrap estimators for ${\bf V}(\widehat{\bm{\theta}}_{\cal L}),{\bf V}(\widehat{\bm{\eta}}_{\cal L}),{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L}),$ and ${\bf V}(\widehat{\bm{\eta}}_{\cal U})$. The regularity conditions for consistent bootstrap variance estimation are outlined in Theorem 3.10 (i) of [41]. We also refer readers to [21], which showed that bootstrap-based variance provides valid but potentially conservative inference.
This result indicates that a greater reduction in variance for the ML-assisted estimator is associated with larger values of ${\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})$ and smaller values of ${\bf V}(\widehat{\bm{\eta}}_{\cal L})$, ${\bf V}(\widehat{\bm{\eta}}_{\cal U})$, and $\rho$. The variance reduction term $[{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})^{\rm T}({\bf V}(\widehat{\bm{\eta}}_{\cal L})+\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U}))^{-1}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})]_{kk}$ can also serve as a metric for selecting the optimal ML model in ML-assisted inference.
Our next result shows that three existing methods, i.e., PPI, PPI++, and PSPA, are asymptotically equivalent to PSPS with different weighting matrices. A broader class for consistent estimator of ${\bm{\theta}}^{*}$ is $\widehat{{\bm{\theta}}}(\bm{\omega})=\bm{\omega}^{\rm T}\widehat{\bm{\eta}}_{{\cal U}}-(\bm{\omega}^{\rm T}\widehat{{\bm{\theta}}}_{{\cal L}}-\widehat{\bm{\eta}}_{{\cal L}})$, where $\bm{\omega}$ is a $K\times K$ matrix. The consistency of $\widehat{{\bm{\theta}}}(\bm{\omega})$ for ${\bm{\theta}}^{*}$ only requires $\bm{\omega}^{\rm T}(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}\mathbf{0}$. Since $(\widehat{\bm{\eta}}_{{\cal U}}-\widehat{\bm{\eta}}_{{\cal L}})\stackrel{{\scriptstyle P}}{{\rightarrow}}\mathbf{0}$, assigning arbitrarily fixed weights for will satisfy the condition. However, the choice of weights influences the efficiency of the estimator as illustrated in proposition 2 later.
@@
Assuming equation (10) and regularity condition for the asymptotic normality of current ML-assisted estimator holds. For any M-estimation problem, we have
$n^{\frac{1}{2}}(\widehat{{\bm{\theta}}}\left({\rm diag}(\bm{\omega}_{\textnormal{ele}}){\bf C}\right)-\widehat{{\bm{\theta}}}_{\textnormal{{PSPA}}})\stackrel{{\scriptstyle D}}{{\rightarrow}}{\bf 0},n^{\frac{1}{2}}(\widehat{{\bm{\theta}}}\left({\rm diag}(\bm{\omega}_{\textnormal{tr}}){\bf C}\right)-\widehat{{\bm{\theta}}}_{\textnormal{{PPI++}}})\stackrel{{\scriptstyle D}}{{\rightarrow}}{\bf 0},n^{\frac{1}{2}}(\widehat{{\bm{\theta}}}\left({\rm diag}({\bf 1}){\bf C}\right)-\widehat{{\bm{\theta}}}_{\textnormal{{PPI}}})\stackrel{{\scriptstyle D}}{{\rightarrow}}{\bf 0}.$
@@
This demonstrates that for M-estimation problems, our method is asymptotically equivalent to PSPA, PPI++, and PPI with the respective weights ${\rm diag}(\bm{\omega}_{\textnormal{ele}}){\bf C}$, ${\rm diag}(\bm{\omega}_{\textnormal{tr}}){\bf C}$, and ${\rm diag}({\bf 1}){\bf C}$. Therefore, PSPS can be viewed as a generalization of these existing methods.
@@
Together with proposition 1, our results demonstrate that our protocol provides a more efficient estimator compared to existing methods for the M-estimation problems. Furthermore, the applicability of our protocol is not limited to M-estimation and only requires summary statistics as input. It also indicates that in a setting of federated data analysis [24] where individual-level data are not available, PSPS proves to be the optimal approach for combining shared summary statistics.
Remark 3.
PPI++ [4] employs a power-tuning scalar for variance reduction in ML-assisted inference. This scalar is obtained by minimizing the trace or possibly other scalarization of the estimator’s variance-covariance matrix. However, the asymptotic variance of PSPS is always equal to or smaller than that of PPI++, irrespective of the scalarization chosen by researchers. This advantage arises because PSPS utilizes a $K\times K$ power tuning matrix, $\bm{\omega}$, for variance reduction, where $K$ represents the dimensionality of parameters. This matrix facilitates information sharing across different parameter coordinates, thereby enhancing estimation precision. The choice of weighting matrix in PSPS also allows for element-wise variance reduction, reducing each diagonal element of the variance-covariance matrix. In contrast, the single scalar in PPI++ can only target overall trace reduction or variance reduction of a specific element. A detailed example is provided in appendix B. Only in one-dimensional parameter estimation tasks, such as mean estimation, PPI++ and PSPS exhibit the same asymptotic variance.
@@ 3.3 Extensions
We also provide several extensions to ensure the broad applicability of our method.
@@ 3.3.1 Labeled data and unlabeled data are not independent
Here, we relax the assumption that the labeled data and unlabeled data are independent. When they are not independent, this can lead to the non-zero covariance between the $\widehat{\bm{\eta}}_{\cal L}$ and $\widehat{\bm{\eta}}_{\cal U}$. Consider a broader class of summary statistics asymptotically satisfying
$n^{1/2}\left(\begin{array}[]{c}\widehat{{\bm{\theta}}}_{\cal L}-{\bm{\theta}}^{*}\\ \widehat{\bm{\eta}}_{\cal L}-\bm{\eta}\\ \widehat{\bm{\eta}}_{\cal U}-\bm{\eta}\\ \end{array}\right)\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}\left\{\left(\begin{array}[]{c}{\bf 0}_{K}\\ {\bf 0}_{K}\\ {\bf 0}_{K}\end{array}\right),\left(\left(\begin{array}[]{ccc}{\bf V}(\widehat{\bm{\theta}}_{\cal L})&{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})\\ {\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})&{\bf V}(\widehat{\bm{\eta}}_{\cal L})&\sqrt{\rho}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})\\ {\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})&\sqrt{\rho}{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U})&\rho{\bf V}(\widehat{\bm{\eta}}_{\cal U})\end{array}\right)\right)\right\}$
We can similarly employ the one-step debiasing $\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}^{\texttt{ no-indep}}=\widehat{\bm{\omega}}_{0}^{\rm T}\widehat{\bm{\eta}}_{\cal U}-\widehat{\bm{\omega}}_{0}(\widehat{\bm{\eta}}_{\cal L}-\widehat{{\bm{\theta}}})$ where $\widehat{\bm{\omega}}_{0}=(\widehat{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})-\widehat{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{\rm T}(\widehat{\bf V}({\widehat{\bm{\eta}}_{\cal L}})+\widehat{\bf V}({\widehat{\bm{\eta}}_{\cal U}})-2\widehat{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{-1}$ and $\widehat{\operatorname{Var}}(\widehat{\bm{\theta}}_{\textnormal{{PSPS}}}^{\texttt{ no-indep}})=\widehat{\operatorname{Var}}(\widehat{\bm{\theta}}_{\cal L})-(\widehat{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})-\widehat{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{\rm T}({\bf V}({\widehat{\bm{\eta}}_{\cal L}})+\widehat{\bf V}({\widehat{\bm{\eta}}_{\cal U}})-2\widehat{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))^{-1}(\widehat{\bf V}(\widehat{\bm{\theta}}_{\cal L},\widehat{\bm{\eta}}_{\cal L})-\widehat{\bf V}(\widehat{\bm{\eta}}_{\cal L},\widehat{\bm{\eta}}_{\cal U}))$. The theoretical guarantees of the proposed estimator can be similarly derived by theorem 1.
@@ 3.3.2 Sensitivity analysis for distributional shift between labeled and unlabeled data
The other assumption of our approach is that the labeled and unlabeled data are identically distributed so that we can ensure $\widehat{\bm{\eta}}_{\cal L}-\widehat{\bm{\eta}}_{\cal U}\stackrel{{\scriptstyle P}}{{\rightarrow}}{\bf 0}$ and validity of PSPS results. To address the potential violation of this assumption, we introduce a sensitivity analysis with hypothesis testing for the null $H_{0}:\bm{\eta}_{{\cal L},k}=\bm{\eta}_{{\cal U},k}$ with test statistics $\frac{\widehat{\bm{\eta}}_{{\cal L},k}-\widehat{\bm{\eta}}_{{\cal U},k}}{\sqrt{\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{{\cal L},k})+\widehat{\operatorname{Var}}(\widehat{\bm{\eta}}_{{\cal U},k})}}\stackrel{{\scriptstyle D}}{{\rightarrow}}\mathcal{N}(0,1)$ to assess if $\bm{\eta}_{{\cal L},k}$ and $\bm{\eta}_{{\cal U},k}$ are significantly different. Here, the subscript $k$ indicates the $k$-th coordinate. We recommend using p-value $<0.1$ as evidence for heterogeneity and caution the interpretation of results from ML-assisted inference.
3.3.3 ML-assisted FDR control